# Indicator Functions Detect Tangentially Transient Behaviour on Decaying Normally Hyperbolic Invariant Manifolds

### v1

@@ 2 Example: The motion of an electrically charged particle around a perturbed magnetic dipole
In the particular example of the system defined by Eq.1 the normal hyperbolicity of the unperturbed NHIM $M^{P}_{0}$ vanishes at one boundary of the cylinder segment and thereby it is clear that any perturbation of the partial integrability of the system triggers the decay of this NHIM $M^{P}_{0}$ surface beginning from this boundary with vanishing normal hyperbolicity. Therefore this dipole example is an ideal system for studying the events occurring during the decay of a NHIM $M_{0}$ under well-controlled conditions. For more details on all these considerations including detailed numerical illustrations see [17, 20] and references therein.
@@ 3 Decay of the NHIM: Numerical Results for the Delay Time and for the Poincaré Map
The delay time is a natural tool to study the phase space of open Hamiltonian systems. In general, the phase space structure indicators are useful to visualize invariant manifolds in the phase space. In particular, they are useful to find KAM islands, unstable and stable manifolds of NHIMs and to obtain information on their bifurcations when we change the parameters in the system. The details of the ideas of phase space structure indicators like the fast Lyapunov indicator, Lagrangian descriptors, Birkoff averages, the classical action, and delay time can be found in the references [21, 22, 23, 24, 25, 26, 20]. For more information on the recent developments in this topic see the references contained in [27, 28].
@@
Where $r_{0}$ is the distance of the initial point to the origin at the initial time, $r^{+}$ the distance at the time $\tau$ of the forward integration, $r^{-}$ the distance at time $\tau$ of the backward integration, and $v^{+}$ and $v^{-}$ are the asymptotic velocities of the particle. For this Hamiltonian system, both velocities coincide $v^{+}=v^{-}$. More details about the delay time as a phase space structure indicator are contained in the reference [20].
@@
Figure 1: Time delay indicator on colour scale for the unperturbed case, $\epsilon=0$ ( see the homoclinic tangle formed by the stable and unstable manifolds of the NHIM $M_{0}$ in Fig.1 in [29] and Fig.2 in [17] for different values of $L_{z}$ ). The integration time $\tau=500$.
To give an interpretation of the figures of the indicator function it is instructive to compare them with the corresponding plots of the inner NHIM $M_{0}$ structure. These plots are constructed by the following method: First, we construct the Poincaré map in the intersection surface $z=0$ where the intersection orientation is irrelevant because of the reflection symmetry of the system in this intersection plane. Second, we restrict this map to the 2-dimensional NHIM $M^{P}_{0}$ surface. Thereby we obtain the Poincaré map for the internal dynamics of the NHIM $M_{0}$, which is a 2-dof dynamics corresponding to a Poincaré map having a 2-dimensional domain. In the following, we call this map of the internal dynamics of the NHIM “the restricted map”.
The Poincaré map of the NHIM $M_{0}$ has a 1:1 projection on the $\phi$$L_{z}$ plane and therefore we show plots of the restricted map by this projection on the $\phi$$L_{z}$ plane. As usual, we present the map graphically by showing many iterates of a moderate number of initial points, for more details see [20]. In Fig.2, we show this Poincaré map for the unperturbed case $\epsilon=0$ and in the various parts of Fig.3 we show the perturbed Poincaré map for the same 4 values of $\epsilon$ which we will use in the various parts of Fig.4.
@@
(b) $\epsilon=0.1$
(c) $\epsilon=0.12$
(d) $\epsilon=0.2$
Figure 3: Poincaré maps of the NHIM $M_{\epsilon}$ for the perturbation parameter values $\epsilon=0.05,0.1,0.12,0.2$
(a) $\epsilon=0.05,\;\phi=-\pi/4$
(b) $\epsilon=0.05,\;\phi=\pi/4$
(c) $\epsilon=0.1,\;\phi=-\pi/4$
(d) $\epsilon=0.1,\;\phi=\pi/4$
(e) $\epsilon=0.12,\;\phi=-\pi/4$
(f) $\epsilon=0.12,\;\phi=\pi/4$
(g) $\epsilon=0.2,\;\phi=-\pi/4$
(h) $\epsilon=0.2,\;\phi=\pi/4$
Figure 4: The time delay function plotted on the $r-L_{z}$ plane for the 4 values $\epsilon=0.05,0.1,0.12,$ and $0.2$ of the perturbation parameter and for the 2 values $\phi=-\pi/4$ and $\phi=\pi/4$ of the azimuth angle. The integration time is $\tau=500$. The value of the time delay is colour coded according to the right hand colour bar.
@@
In Fig.3 (a) we also observe how the decay starts from small values of $L_{z}$. However, the decaying region is still separated from the separatrix structure mentioned above by primary KAM curves in between. For the internal dynamics restricted to the NHIM $M_{0}$ and therefore also for the restricted Poincaré map the KAM curves are impenetrable. Accordingly, the separatrix structure still is a truly invariant remaining part of the NHIM $M_{\epsilon}$ for such small perturbations.
The situation changes drastically for larger values of the perturbation. In Fig.3 (b) the KAM lines below the large secondary structure have already been broken and thereby the former separatrix structure has established contact with the decay front of the NHIM $M_{\epsilon}$. Then strictly speaking the complete separatrix structure is no longer a remaining invariant part of the NHIM $M_{\epsilon}$. Iterates of initial points from this separatrix region at first move chaotically inside of the separatrix structure. However after a sufficient number of iterations, they come close to the decay front, possibly cross it and leave the surviving NHIM $M_{\epsilon}$ surface tangentially.
To illustrate this effect pictorially several initial points on the line $\phi=-\pi/4$ have been chosen and their iterates are also included in the figure as long as these iterated points could be stabilized on the former NHIM $M^{P}_{\epsilon}$ surface by our method to construct the restricted map. See the irregularly scattered points in the former separatrix region and below the remaining large KAM islands. These trajectories have a transient existence on the remaining parts of the NHIM $M^{P}_{\epsilon}$ surface only.
@@
This property can be assured along the following considerations. We have already mentioned that the NHIM surface projects 1:1 into the $\phi-L_{z}$ plane. And on the NHIM $M_{\epsilon}$ the coordinates $r$ and $p_{r}$ are always close to the values $r=1$ and $p_{r}=0$. Therefore the $r-L_{z}$ plane at $p_{r}=0$ and at any value of $\phi$ should work well. With the Fig.3 in mind, we have chosen the two values $\phi=\pm\pi/4$ with the intention that the plot at one value should be heavily influenced by the big secondary islands and the other plot should emphasise the separatrix structure.
@@
Globally $W^{s}(M_{\epsilon})$ and $W^{u}(M_{\epsilon})$ form homoclinic intersections and build up a chaotic tangle. Therefore the stable and unstable manifold form a fractal of tendrils. In the end, we have in $S$ a fractal collection of 1-dimensional intersection lines with $W^{s}(M_{\epsilon})$ leading to a fractal collection of lines of singularities of the time delay function. This is exactly what we see in Fig.1. Between the curves of singularities, the time delay drops to finite values. However in regions of a high density of lines of singularities with only very small gaps in between them also inside of the gaps we have rather high finite values of the time delay. In the plots, the whole region has obtained yellow colour. However, in a comparison between Figs.1 and 5, we see that under magnification we can resolve more gaps in the fractal intersection structure.
Next, we consider nonzero perturbations. We know already from the plots of the restricted map in Fig.3 that along $\phi=-\pi/4$ there is a lot of decay on the NHIM $M^{P}_{\epsilon}$ whereas along $\phi=\pi/4$ the NHIM $M^{P}_{\epsilon}$ forms a large island and thereby the NHIM $M^{P}_{\epsilon}$ surface survives. Accordingly on the indicator function in the plane along $\phi=\pi/4$ sharp fractal structures remain, even though they change in some details, whereas in the plane along $\phi=-\pi/4$ a considerable part of the former fractal structures is converted into a very diffuse region of still large but not infinite values of the time delay.
@@
In the final part of this section, we point out some details of the delay plots which demonstrate how closely the indicator function follows the changes in the NHIM $M_{\epsilon}$ structure. In Fig.4 we also observe surviving clearly fractal structures for high values of $L_{z}$ corresponding to the surviving parts of the NHIM $M^{P}_{\epsilon}$ at high values of $L_{z}$ also for $\phi$ around $-\pi/4$. This parallelism explains in which way indicator functions can illustrate the decay scenario of NHIMs.
In Fig.1 we observe very high values of the time delay around $\phi=1.2$, $L_{z}=0.9$ representing a fractal of singularities which is not well resolved. Fig.5 shows that under higher resolution and a related rescaling of the colour bar, the plot still represents a fractal of singularities. The Figs. 4(a) and 4(b) show that these very high values of the time delay survive the small perturbation of $\epsilon=0.05$. The structures are only deformed and shifted a little. These structures in the delay functions are caused by initial conditions converging towards the separatrix structure along the part of $W^{s}(M_{\epsilon})$ lying over the separatrix structure. Remember that according to the foliation theorem ( chapter 5 of [1] ) $W^{s}(M_{\epsilon})$ transports the inner structure of the NHIM $M_{\epsilon}$ along it.
Next we increase the perturbation to the value $\epsilon=0.1$ in Figs.4(c) and 4(d). Here the very high value of the time delay for $\phi\approx 1.2$ and $L_{z}\approx 0.9$ is no longer present. The better resolution under the magnification in Fig.6 shows that for this higher perturbation, the true singularities pointed out above are gone. We also observe that for $\phi=-\pi/4$ in Fig.4(c) the background value has dropped considerably whereas for $\phi=\pi/4$ in Fig.4(d) at least some higher background can be recognised. In this form, we get a clue that for smaller perturbations there has been some important structure present which becomes destroyed for increasing perturbation.
Figure 5: Magnification of the time delay indicator from Fig.1 on colour scale for the unperturbed case, $\epsilon=0$. The integration time $\tau=500$.
Figure 6: Magnification of time delay indicator from Fig.4 (c) at $\epsilon=0.1$, $\phi=-\pi/4$. The integration time $\tau=500$.
Finally, we point out some analogies between the indicator function and the Poincaré map in more detail. In Fig.2(a) we see that for $\epsilon=0.05$ the large separatrix structure is still confined on both sides by primary KAM curves. It has not yet established any contact with the decay front. Therefore trajectories in this separatrix structure can not escape by tangential motion on the surviving parts of the NHIM $M_{\epsilon}$. This statement is valid for all values of $\phi$.
@@
By looking at Fig.3(b) it becomes understandable that initial conditions in the former separatrix region around $\phi=\pi/4$ and above the large secondary island needs a longer time to diffuse tangentially to the decaying front than the ones starting around $\phi=-\pi/4$. Accordingly, there still exist transient remnants of the stable manifold of the former separatrix structure near $\phi=\pi/4$ which also have influence on the indicator function when they intersect its domain in contrast to the behaviour around $\phi=-\pi/4$.
This explains why in Fig.4(d) we still observe long transients near $r=1.1,L_{z}=0.9$ in contrast to Fig.4(c) where we observe rather short transients only. A comparison between the whole scenarios of Fig.4 and Fig.3 makes it easy to understand why in total the transient effects fade out with increasing perturbation.

### v2

@@ 2 Example: The motion of an electrically charged particle around a perturbed magnetic dipole
In this particular example, the normal hyperbolicity of the unperturbed NHIM $M^{P}_{0}$ vanishes at one boundary of the cylinder segment and thereby it is clear that any perturbation of the partial integrability of the system triggers the decay of this NHIM $M^{P}_{0}$ surface beginning from this boundary with vanishing normal hyperbolicity. Therefore, this dipole example is an ideal system for studying the events occurring during the decay of a NHIM $M_{\epsilon}$ under well-controlled conditions. For more details on all these considerations, including detailed numerical illustrations, see [18, 21] and references therein.
3 The Numerical Algorithm to Follow a NHIM Along a Curve of the Perturbation Parameter
We give a short description of our method to search for NHIMs and for the pictorial presentation of their internal dynamics. And we include comments which may help to clarify our method.
The method begins with a search for the local branch of the stable manifold of an unstable invariant subset of codimension 2 in the Poincaré map. This stable manifold has codimension 1 and therefore we expect that a curve of dimension 1 placed into an appropriate region of the domain of the map intersects this stable manifold in isolated points. We search for one of such intersection points.
For the starting case, $\epsilon=0$, we know the NHIM and its position. Let us call it $M_{0}$. And now we search for the NHIM for some small value of $\epsilon\neq 0$. Let us call it $M_{\epsilon}$. Because of continuity and persistence, $M_{\epsilon}$ is located near $M_{0}$. So we select some tubular neighbourhood $U$ around $M_{0}$ such that we are confident that $M_{\epsilon}$ lies within $U$. Next, we place some curve $C$ into $U$ such that we suppose that $C$ intersects the local branch of $W^{u}(M_{\epsilon})$. We place many initial points along $C$ and calculate for each one of these points the number of iterations of the full 4-dimensional Poincaré map which lie still within $U$.
If one of these points would lie exactly on the local branch of the stable manifold of an invariant subset, then this number of iterations within $U$ would be infinite. Of course, numerically we only detect the region of $C$ where this number becomes maximal. Then, we zoom in as fine as our numerical precision allows to get this maximum with the best possible precision. The corresponding trajectory ( sequence of iterated points ) first approaches the invariant subset along its stable manifold, stays close to the subset for a while and finally leaves the neighbourhood of the invariant subset again along its unstable manifold. Now, we truncate this trajectory sufficiently long before it leaves $U$ such that we are confident that the last point we keep is still very close to the invariant subset.
We use this last point which we keep as our new initial point. We place an appropriate new curve $C$ through this last point. This time $C$ can be rather short. And we search for the point along the new curve $C$ which keeps the trajectory inside of $U$ for the longest possible sequence of iterated points. And we continue this procedure. Thereby we continue the sequence of points in the close neighbourhood of $M_{\epsilon}$ until we have a long sequence of points very close to the NHIM we are looking for.
So the procedure consists of iterated applications of the full 4-dimensional map with repeated interruptions to perform the fine-tuning along some curve $C$. Many times it is convenient to place the curve $C$ into the direction of the momentum normal to the NHIM surface. This fine-tuning is a version of the control of chaos where we keep some numerical trajectory on the stable manifold of some invariant subset, see also [22, 23]
We must understand why this trajectory represents some substructure of the internal dynamics of the NHIM. The answer is given by the foliation theorem of stable and unstable manifolds of NHIMs, see chapter 5 in [1]. Basically ( expressed in a somewhat oversimplified manner ) it says that these stable and unstable manifolds are the Cartesian product of a line with the internal structure of the NHIM. Then our initial point belongs to some internal structure of the NHIM and the whole trajectory stays on this particular internal structure and represents it. Behind this foliation property stands the property of the full dimensional Poincaré map in the neighbourhood of the NHIM to have a natural factorization into a normal factor and a tangential factor. Our numerical procedure filters out automatically the tangential factor.
For some well-chosen collection of initial curves $C$ placed over the various important internal structures of the NHIM, we can produce a corresponding collection of trajectories on the NHIM which represent all important internal structures of the NHIM. To this end, we can try out many initial curves $C$ and select the ones which lead to a good pictorial presentation of the NHIM structures, including important KAM islands, important secondary structures, chaos stripes, etc, in analogy to what we are used to do when constructing plots of 2-dimensional Poincaré maps. By this procedure, we obtain simultaneously the position of the NHIM and a pictorial representation of its internal dynamics. The plots we show are the projection of the 2-dimensional NHIM surface into some coordinate plane of the domain of the 4-dimensional Poincaré map. Preferably, we use a coordinate plane giving a 1:1 projection, whenever this is possible.
Now assume that we have the NHIM for some value $\epsilon_{1}$ of the perturbation and want to construct it for some new value $\epsilon_{2}$ which is a little larger. Then we place a new neighbourhood $U$ around the $M_{\epsilon_{1}}$ and apply the same idea as before to construct the NHIM for the new larger perturbation. So we proceed in sufficiently small steps towards larger perturbation. At some point, we come to a value of $\epsilon$ where the NHIM starts to lose normal hyperbolicity at least locally at some point.
Here the reader may ask how we notice the loss of normal hyperbolicity and how we can check normal hyperbolicity. First, the whole search strategy relies heavily on the existence of a stable manifold of codimension 1, and this, in turn, relies on the normal hyperbolicity. In this sense, we can interpret the success of our method as a numerical indication of the presence of normal hyperbolicity at least in the local region of the NHIM where we are looking for some internal NHIM structure. There are many examples of the success of this approach, see e.g Fig.7 in [5], Figs.7 and 8 in [24], Figs.12 and 13 in [25], Figs.16, 17, 18, and 19 in [26], Figs.14 and 15 in [27] and Fig.13 in [28] and several more examples which are not yet published.
We can do even more to check normal hyperbolicity directly in some regions of the NHIM. If we know some periodic orbit within the NHIM region, then we can calculate the monodromy matrix of this orbit and obtain its stability properties in the tangential and in the normal direction of the NHIM. Then we see at which value of some parameter the tangential instability of the corresponding fixed point in the reduced map becomes larger than the normal instability, i.e. at which value it loses normal hyperbolicity. A nice example of this check has been shown in [5]. Here at some parameter value, an important fixed point on the NHIM loses normal hyperbolicity. Exactly at this parameter value the NHIM surface grows a singularity and breaks and the numerical search method starts to fail in the neighbourhood of this fixed point, the NHIM surface grows a hole. In [28] another example of the creation of a hole in the NHIM and the corresponding breakdown of our search strategy has been presented. Again it happens exactly at the parameter value at which an important periodic orbit inside the NHIM loses its normal hyperbolicity. The formation of the hole in the NHIM surface starts exactly at the position of the fixed point which loses normal hyperbolicity, compare Figs.6 and 13 in [28].
Now it is time to make a few clarifying remarks on our search strategy. It is based only on the very elementary qualitative properties of NHIMs and the trajectories in their neighbourhoods. It only needs the numerical construction of the full 4-dimensional Poincaré map of the dynamics or equivalently it only needs the numerical construction of trajectories in the full 6-dimensional phase space of the system. It never needs any reduction of the dynamics to some centre manifold or some other lower dimensional subsets nor does it need any construction of normal forms or any other analytical approximations. The above mentioned factorization of the dynamics does the job all by itself. The reduction of the dynamics to the centre manifold is an alternative strategy which does this factorization by analytical approximations and by the construction of normal forms, see [29, 30].
As mentioned, this simple method is based on qualitative properties of NHIMs and the trajectories in their neighbourhoods only. Therefore, the main mathematical condition which ever enters the whole procedure is the maximization of the time that the trajectory spends in the neighbourhood of the NHIM, as we change the initial conditions. The whole procedure is numerical and it only relies on the general qualitative properties of NHIMs. Of course, this has the disadvantage that we can not derive the conclusions and explanations of the dynamics which one might draw from reductions and normal forms [31, 32, 33] or the parametrization method [34]. However, our method can give us important details about the rich bifurcation process regardless of any hypothesis considered in analytical methods.
Finally, we proceed to some remarks on the events which happen as soon as the decay of the NHIM has started. Suppose that a chaos stripe comes into contact with the decay boundary of the NHIM and we start some trajectory over the part of this chaos strip which still lies behind the boundary in the part of the NHIM which still shows normal hyperbolicity. Then we obtain an initial part of the trajectory belonging to the chosen initial point by our search method. However, after some number of iterations, this trajectory crosses the decay boundary and enters a region where normal hyperbolicity is already lost. Then also the stable manifold as a codimension 1 surface is lost. In the best case a collection of lower-dimensional fragments still exist.
Then our search method no longer finds a clearly peaked maximum of the number of iterations along the curve $C$. It only finds a region of high values with several relative maxima and minima, as a typical example see Fig.4 in [19]. Then the intent to zoom in on some true infinity fails and the method breaks down. We interpret the initial segment of the trajectory up to the breakdown of the search method as a transient trajectory on the NHIM. In our example of the perturbed magnetic dipole such chaos regions with contact with the decay boundary can have large areas and accordingly, they have a large impact on plots of indicator functions. This is the topic of the present article.
@@ 4 Decay of the NHIM: Numerical Results for the Delay Time and for the Poincaré Map
The delay time is a natural tool to study the phase space of open Hamiltonian systems. In general, the phase space structure indicators are useful to visualize invariant manifolds in the phase space. In particular, they are useful to find KAM islands, unstable and stable manifolds of NHIMs and to obtain information on their bifurcations when we change the parameters in the system. The details of the ideas of phase space structure indicators like the fast Lyapunov indicator, Lagrangian descriptors, Birkhoff averages, the classical action, and delay time can be found in the references [21, 35, 36, 37, 38, 39, 40]. For more information on the recent developments in this topic, see the references contained in [41, 42].
@@
Here $r_{0}$ is the distance of the initial point to the origin at the initial time, $r^{+}$ the distance at the time $\tau$ of the forward integration, $r^{-}$ the distance at time $\tau$ of the backward integration, and $v^{+}$ and $v^{-}$ are the asymptotic velocities of the particle. For this Hamiltonian system, both velocities coincide $v^{+}=v^{-}$. More details about the delay time as a phase space structure indicator are contained in the reference [21].
@@
Figure 1: Time delay indicator on a colour scale for the unperturbed case, $\epsilon=0$ ( see the homoclinic tangle formed by the stable and unstable manifolds of the NHIM $M_{0}$ in Fig.1 in [43] and Fig.2 in [18] for different values of $L_{z}$ ). The integration time $\tau=50$.
To give an interpretation of the figures of the indicator function, it is instructive to compare them with the corresponding plots of the inner NHIM $M_{0}$ structure. These plots are constructed by the following method: First, we construct the Poincaré map in the intersection surface $z=0$ where the intersection orientation is irrelevant because of the reflection symmetry of the system in this intersection plane. Second, we restrict this map to the 2-dimensional NHIM $M^{P}_{0}$ surface. Thereby, we obtain the Poincaré map for the internal dynamics of the NHIM $M_{0}$, which is a 2-dof dynamics corresponding to a Poincaré map having a 2-dimensional domain. In the following, we call this map of the internal dynamics of the NHIM “the restricted map”.
The Poincaré map of the NHIM $M_{0}$ has a 1:1 projection on the $\phi$$L_{z}$ plane and therefore we show plots of the restricted map by this projection on the $\phi$$L_{z}$ plane. As usual, we present the map graphically by showing many iterates of a moderate number of initial points, for more details see [21]. In Fig.2, we show this Poincaré map for the unperturbed case $\epsilon=0$ and in the various parts on the left column of Fig.3 we show the perturbed Poincaré map for 4 values of $\epsilon$.
@@
(b) $\epsilon=0.05,\;\phi=-\pi/4$
(c) $\epsilon=0.05,\;\phi=\pi/4$
(d) $\epsilon=0.1$
(e) $\epsilon=0.1,\;\phi=-\pi/4$
(f) $\epsilon=0.1,\;\phi=\pi/4$
(g) $\epsilon=0.12$
(h) $\epsilon=0.12,\;\phi=-\pi/4$
(i) $\epsilon=0.12,\;\phi=\pi/4$
(j) $\epsilon=0.2$
(k) $\epsilon=0.2,\;\phi=-\pi/4$
(l) $\epsilon=0.2,\;\phi=\pi/4$
Figure 3: On the left column, the Poincaré maps of the NHIM $M_{\epsilon}$ for the perturbation parameter values $\epsilon=0.05,0.1,0.12,0.2$. On the central and right column the time delay function plotted on the $r-L_{z}$ plane for the 4 values $\epsilon=0.05,0.1,0.12,$ and $0.2$ of the perturbation parameter and for the 2 values $\phi=-\pi/4$ and $\phi=\pi/4$ of the azimuth angle. The integration time is $\tau=50$. The value of the time delay is colour-coded according to the right-hand colour bar.
@@
In Fig.3 (a) we also observe how the decay starts from small values of $L_{z}$. However, the decaying region is still separated from the separatrix structure mentioned above by primary KAM curves in between. For the internal dynamics restricted to the NHIM $M_{0}$ and therefore also for the restricted Poincaré map, the KAM curves are impenetrable. Accordingly, the separatrix structure still is a truly invariant remaining part of the NHIM $M_{\epsilon}$ for such small perturbations.
The situation changes drastically for larger values of the perturbation. In Fig.3 (d) the KAM lines below the large secondary structure have already been broken and thereby the former separatrix structure has established contact with the decay front of the NHIM $M_{\epsilon}$. Then strictly speaking the complete separatrix structure is no longer a remaining invariant part of the NHIM $M_{\epsilon}$. Iterates of initial points from this separatrix region at first move chaotically inside of the separatrix structure. However after a sufficient number of iterations, they come close to the decay front, possibly cross it and leave the surviving NHIM $M_{\epsilon}$ surface tangentially.
To illustrate this effect pictorially, several initial points on the line $\phi=-\pi/4$ have been chosen and their iterates are also included in the figure as long as these iterated points could be stabilized on the former NHIM $M^{P}_{\epsilon}$ surface by our method to construct the restricted map. See the irregularly scattered points in the former separatrix region and below the remaining large KAM islands. These trajectories have a transient existence on the remaining parts of the NHIM $M^{P}_{\epsilon}$ surface only.
@@
This property can be assured along the following considerations. We have already mentioned that the NHIM surface projects 1:1 into the $\phi-L_{z}$ plane. And on the NHIM $M_{\epsilon}$ the coordinates $r$ and $p_{r}$ are always close to the values $r=1$ and $p_{r}=0$. Therefore the $r-L_{z}$ plane at $p_{r}=0$ and at any value of $\phi$ should work well. With the left column of Fig.3 in mind, we have chosen the two values $\phi=\pm\pi/4$ with the intention that the plot at one value should be heavily influenced by the big secondary islands and the other plot should emphasise the separatrix structure.
@@
Globally $W^{s}(M_{\epsilon})$ and $W^{u}(M_{\epsilon})$ form homoclinic intersections and build up a chaotic tangle. Therefore the stable and unstable manifold form a fractal of tendrils. In the end, we have in $S$ a fractal collection of 1-dimensional intersection lines with $W^{s}(M_{\epsilon})$ leading to a fractal collection of lines of singularities of the time delay function. This is exactly what we see in Fig.1. Between the curves of singularities, the time delay drops to finite values. However, in regions of a high density of lines of singularities with only very small gaps in between them, also inside of the gaps we have rather high finite values of the time delay. In the plots, the whole region has obtained yellow colour. However, in a comparison between Figs.1 and 4, we see that under magnification we can resolve more gaps in the fractal intersection structure.
Next, we consider nonzero perturbations. We know already from the plots of the restricted map in Fig.3 that along $\phi=-\pi/4$ there is a lot of decay on the NHIM $M^{P}_{\epsilon}$ whereas along $\phi=\pi/4$ the NHIM $M^{P}_{\epsilon}$ forms a large island and thereby the NHIM $M^{P}_{\epsilon}$ surface survives. Accordingly, on the indicator function in the plane along $\phi=\pi/4$ sharp fractal structures remain, even though they change in some details, whereas in the plane along $\phi=-\pi/4$ a considerable part of the former fractal structures is converted into a very diffuse region of still large but not infinite values of the time delay.
@@
In the final part of this section, we point out some details of the delay plots which demonstrate how closely the indicator function follows the changes in the NHIM $M_{\epsilon}$ structure. In Fig.3 we observe surviving fractal structures for high values of $L_{z}$ corresponding to the surviving parts of the NHIM $M^{P}_{\epsilon}$ at high values of $L_{z}$ also for $\phi$ around $-\pi/4$. This parallelism explains in which way indicator functions can illustrate the decay scenario of NHIMs.
In Fig.1 we observe very high values of the time delay around $\phi=1.2$, $L_{z}=0.9$ representing a fractal of singularities which is not well resolved. Fig.4 shows that under higher resolution and a related rescaling of the colour bar, the plot still represents a fractal of singularities. The Figs. 3(b) and 3(c) show that these very high values of the time delay survive the small perturbation of $\epsilon=0.05$. The structures are only deformed and shifted a little. These structures in the delay functions are caused by initial conditions converging towards the separatrix structure along the part of $W^{s}(M_{\epsilon})$ lying over the separatrix structure. Remember that according to the foliation theorem ( chapter 5 of [1] ) $W^{s}(M_{\epsilon})$ transports the inner structure of the NHIM $M_{\epsilon}$ along it.
Next we increase the perturbation to the value $\epsilon=0.1$ in Figs.3(e) and 3(f). Here the very high value of the time delay for $\phi\approx 1.2$ and $L_{z}\approx 0.9$ is no longer present. The better resolution under the magnification in Fig.5 shows that for this higher perturbation, the true singularities pointed out above are gone. We also observe that for $\phi=-\pi/4$ in Fig.3(e) the background value has dropped considerably whereas for $\phi=\pi/4$ in Fig.3(f) at least some higher background can be recognised. In this form, we get a clue that for smaller perturbations there has been some important structure present which becomes destroyed for increasing perturbation.
Figure 4: Magnification of the time delay indicator from Fig.1 on colour scale for the unperturbed case, $\epsilon=0$. The integration time $\tau=50$.
Figure 5: Magnification of time delay indicator from Fig.3 (e) at $\epsilon=0.1$, $\phi=-\pi/4$. The integration time $\tau=50$.
Finally, we point out some analogies between the indicator function and the Poincaré map in more detail. In Fig.3(a) we see that for $\epsilon=0.05$ the large separatrix structure is still confined on both sides by primary KAM curves. It has not yet established any contact with the decay front. Therefore, trajectories in this separatrix structure cannot escape by tangential motion on the surviving parts of the NHIM $M_{\epsilon}$. This statement is valid for all values of $\phi$.
@@
By looking at Fig.3(d) it becomes understandable that initial conditions in the former separatrix region around $\phi=\pi/4$ and above the large secondary island need a longer time to diffuse tangentially to the decaying front than the ones starting around $\phi=-\pi/4$. Accordingly, there still exist transient remnants of the stable manifold of the former separatrix structure near $\phi=\pi/4$, which also have an influence on the indicator function when they intersect its domain, in contrast to the behaviour around $\phi=-\pi/4$.
This explains why in Fig.3(f) we still observe long transients near $r=1.1,L_{z}=0.9$ in contrast to Fig.3(e) where we observe rather short transients only. A comparison between the whole scenarios of Fig.3 makes it easy to understand why in total the transient effects fade out with increasing perturbation.